What is the volume of a sample of ethane at 467k and 1.1atm ,if it occupies 405ml at 298k and 1.1atm



1
Expert's answer
2022-03-04T03:42:01-0500

In the question given that,

P1= 1.1 atm, T1= 467 K, V1=?

P2= 1.1 atm , T2= 298 K, V2 = 405 ml


"\frac{P_1V_1}{T_1}= \frac{P_2V_2}{T_2}"


"\frac{1.1V_1}{467}= \frac{1.1×405}{298}"


"V_1= 634.67" ml

So, volume of sample is 634.67 ml